an earth satellite,in circular orbit at an altitude h of 230 km above the earth's surface,has a period T of 89 minutes.What mass of the earth follows from these data?
1
Expert's answer
2019-06-13T09:39:31-0400

Newton's second law:

"\frac{GMm}{(R+h)^2} = ma\\
a= \frac{GM}{(R+h)^2}\\
a= \frac{v^2}{(R+h)}\\
v=\frac{2\pi (R+h)}{T}\\
a=\frac{4 \pi^2 (R+h)}{T^2}\\
\frac{4 \pi^2 (R+h)}{T^2}=\frac{GM}{(R+h)^2}\\"

and then the mass of Earth is

"M=\frac{4 \pi^2 (R+h)^3}{GT^2}=\\
=\frac{4\pi^2 ((6371+230)\cdot 10^3)^3}{6,67\cdot 10^{-11}\cdot (89\cdot 60)^2} = 5.97\cdot 10^{24} kg"